OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-9, 9) and is perpendicular to -9x – 2y = 10

OpenStudy (saifoo.khan):

-9/2

OpenStudy (anonymous):

If the line is perpendicular then it has a negative reciprocal slope. So the slope would be 2/9 (solve the above equation you get y=(-9/2)x-5) Then you have a point. Use point slope form. \[y-y_0=m(x-x_0) \rightarrow y-9=(2/9)(x-(-9))\] Then solve. You should get: y=(2/9)x+11

OpenStudy (anonymous):

I got it mal...

OpenStudy (radar):

slope for the given line -9x-2y=10 when placed in slope intercept form will be -2y=9x+10 becomes y=-(9/2)x-5 The coefficient of x is the slope and it is -(9/2) Any line that is perpendicular to that line of slope -(9/2) will be the negative reciprocal. that means the slope for the perpendicular line will be a (2/9) note that it is positive. and is the reciprocal. Now this line is said to go through P(-9,9). With the slope now known, and the point given we can write the equation for the perpendicular line in the slope intercept form. y-9=m(x--9) that is actually y-9=m(x +9) substituting the value of the slope y-9=(2/9)x+(2/9)9 y=2x/9 + 2 +9 or y=(2/9)x + 11 just as malevolence 19 worked it out.

OpenStudy (anonymous):

Good :D